1.) A cello string vibrates in its fundamental
mode with a frequency of 179 1/s. The vibrating segment is 29.2 cm long and has a
mass of 0.75 g.
Find the tension in the string.
Answer in units of N

2).Find the frequency of the string when it
vibrates in seven segments.
Answer in units of 1/s.
1
Expert's answer
2020-02-20T09:25:21-0500

1)

"v=f\lambda=2lf=\sqrt{\frac{Tl}{m}}"

"2(179)(0.292)=\sqrt{\frac{(0.292)T}{0.00075}}"

"T=28.1\ N"

2)


"f_7=7f=7(179)=1253\ s^{-1}"